The number of 10 centavo coins in a box was equal to the number of 25 centavo coins. There were five times as many 5 centavo coins as 25 centavo coins. All these coins togethr totaled P120. How many 5 centavo coins were there in a Box?
Answer by mananth(16946)

The number of 10 centavo coins =x
number of 25 centavo coins =x
five times as many 5 centavo coins as 25 centavo coins. = 5x
total= P120.
...
10x+25x+5*5x=120
60x=120
/60
x=2
5 centavo coins = 5x = 5*2 = 10
...
m.ananth@hotmail.ca
10*2+25*2+5*10=120